The density of carbon dioxide gas was measured at pressure of 0.8 atm and temperature of 37 c. it was found to be equal to:
1
Expert's answer
2016-10-26T13:01:08-0400
Assume that carbon dioxide here behaves as an ideal gas and we can use the ideal gas law:

PV = nRT

Let us derive density in [mol/l]:

D = n/V = P/RT

And then in g/l:

D = m/V = nM/V = PM/RT

Here R = 8.314 (L*kPa)/(mol*K)
P = 0.8 atm * 101.3 kPa/atm =
T = 310 K
M = 44 g/mol

D = 81.04 * 44 / (8.314 * 310) = 1.38 g/L or 1.38 * 10^(-3) g/cm^3